OpenStudy (anonymous):

Write the equation of the line that passes through point p and has the given slope? @ehuman P (10, 1) slope 1/5

OpenStudy (skyz):

Its to early fo math (ಥ_ಥ) xDD MMk i had to say that :P

OpenStudy (skyz):

Im never good with graphs or slopes.

OpenStudy (akashdeepdeb):

P (10, 1) slope 1/5 Do you know the point slope form a of a line? Slope = m = Rise/Run = \(\frac{\delta y}{\delta x}\) = \(\frac{change~~in~~y}{change~~in~~x}\) = m = \(\frac{y-y_1}{x-x_1}\) Or \(y-y_1 = m(x-x_1)\) And you are given a point through which the line passes = P(10,1) = (\(x_1,y_1\)) Slope (m) = \(\frac{1}{5}\) Put them in the formula and you'll get the line equation. :) Understood? :)

OpenStudy (anonymous):

@AkashdeepDeb So far I understand that ?-1=1/5 (?-10)

OpenStudy (akashdeepdeb):

?-1=1/5 (?-10) Better to write it like this: y-1 = 1/5 * (x-10) THAT IS IT. NOW JUST CALCULATE! :D 5*(y-1) = (x-10) 5y - 5 = x-10 x - 5y = 5 Got it? :)